OpenStudy (anonymous):

A motorcyclist changes the velocity of his bike from 20.0 meters/second to 35.0 meters/second under a constant acceleration of 4.00 meters/second2. How long does it take the bike to reach the final velocity?

OpenStudy (silverlark2):

we know that acceleration is the rate of change in velocity, hence a= (v-u)/t (note that v= final velocity and u=intial velocity). Hence, making t the subject of the eqn, we will have t= (v-u)/a . Sub in the values given in the question into the eqn: v= 35m/s , u= 20m/s and a = 4m/s^2 and solve for t.

OpenStudy (anonymous):

Initial Velocity of the motorcyclist=20.0m/s (u) Final Velocity of the motorcyclist=35.0m/s (v) acceleration=4.00m/s^2 According to the first law of motion- v=u +at now making t the subject of the formula- v=u +at (v -u)= at (v-u)/a=t Now substituting the values in the formula- (35 -20)/4 =t 15/4=t 3.75s =t Ans- The Bike took 3.5s to reach the final velocity.